OpenStudy (anonymous):

Find 3 consecutive even numbers where the product of the smaller two numbers is 64 less than the square of the largest number.

OpenStudy (anonymous):

Let the nos. be n,n+2,n+4 \[n \left( n+2 \right)=\left( n+4 \right)^{2}-64\] solve,you get the answer